# Verifiable Natural Language to Linear Temporal Logic Translation: A Benchmark Dataset and Evaluation Suite

### v1

@@ 3.3 Metrics
Each dataset entry includes three paired artifacts:
$\bigl{\{}\,\underbrace{\mathrm{sentence},\,\mathrm{lifted\ sentence\}}}_{\text{NL (grounded \& lifted)}},\;\underbrace{\varphi_{G},\,\varphi_{L}}_{\text{LTL(grounded \& lifted)}},\;\underbrace{\sigma_{good}\models\varphi_{G},\,\sigma_{bad}\not\models\varphi_{G}}_{\text{Traces (holds \& $\neg$ holds)}}\bigr{\}}.$
This rich annotation supports four independent evaluation axes, displayed in Figure 2:
  1. 1.
    Lifting accuracy: The ability to correctly extract substrings from a natural language specification that map to a corresponding atomic proposition in the linear temporal logic.
  2. 2.
    Translation accuracy: The fidelity of mapping grounded or lifted NL to the correct temporal-logic formula.
  3. 3.
    Grounding accuracy: The ability to align logical atomic proposition labels with their corresponding world state condition.
  4. 4.
    Verification accuracy: The capacity to decide satisfaction or violation of a grounded LTL formula on a given trace or traces.
@@ 3.4 Datasets
To use our data synthesis pipeline, we provide three scenario definitions (listed in Appendix A.5) accompanied by action and target references. In total, our synthesis pipeline yields three new datasets, usable for training and evaluation. Though the same generation pipeline is used to produce all three datasets, we note that the input and In this section, we describe each of the included datasets.
@@
Our Warehouse dataset simulates a realistic warehouse retrieval scenario, explicitly designed for scalability and complexity in grounding tasks. Warehouse is our most distinct dataset with its inclusion of all 80 COCO [17] object classes, significantly enriching the semantic and linguistic complexity and variation of atomic propositions (APs). Each entry includes LTL formulas with explicit grounding and alignment at token-level granularity. The dataset also uniquely provides verified positive ("good") and negative ("bad") execution traces for robust validation.
@@
  • Grounded LTL Formula: globally( deliver(bench, loading_dock) or idle or search(wine_glass))
  • APs: prop_1 = “drop off long chair to loading dock”, prop_2 = “idle”, prop_3 = “look for glass for alcoholic beverage”
Search and Rescue.
The Search-and-Rescue (S&R) scenario presents a detailed context of emergency response and robotic aid. It emphasizes multi-step TL relationships, requiring precise event sequencing and robust conditional propositions. Like the Warehouse scenario, it provides complete grounding information and includes scenario-specific LTL templates with both positive and negative execution traces for rigorous verification. The S&R scenario deliberately includes an "avoid" action for the purpose of stress-testing the a framework’s ability to distinguish between APs and LTL operators, because avoid is commonly used to indicate a "globally (negation $\varphi$)" pattern.
Example:
  • Sentence: “If deliver aid to the safe rescuer holds and request assistance holds next, then stay away from the active gas leak holds in the step after that.”
  • Lifted Sentence: “If prop_1 holds and prop_2 holds next, then prop_3 holds in the step after that.”
  • Grounded LTL Formula: globally(deliver_aid(safe_rescuer) and next get_help implies next next avoid(active_gas_leak))
  • APs: prop_1 = deliver aid to safe rescuer, prop_2 = request assistance, prop_3 = avoid active gas leak
Traffic Light Control.
The Traffic Light Control scenario introduces LTL propositions relevant to urban navigation and infrastructure management. This scenario tests an NL-to-LTL system’s ability to accurately ground complex real-world relationships, such as spatial landmarks and persistent behaviors like "infinitely often." Each dataset entry is provided with explicit good and bad traces, detailed proposition grounding, and is constructed via a flexible scenario template, allowing straightforward extension to additional real-world tasks.
Example:
  • Sentence: “Take a photo of the motorcycle on northeast 10th avenue must keep holding until, from some point on, change east light red holds infinitely often.”
  • Lifted Sentence: “prop_1 must keep holding until prop_2 holds infinitely often.”
  • Grounded LTL Formula: photo(motorcycle,northeast_10th_avenue) until globally(finally change(light_east,red))
  • APs: prop_1 = take a photo of motorcycle, northeast 10th avenue, prop_2 = change east light red
@@ 4 Experimental Results
In this section, we present the results of multiple evaluations of NL-to-LTL translation frameworks and components. In Section 4.1, we measure the performance of common natural language lifting approaches, evaluated on four existing datasets in addition to the three datasets we present in VLTL-Bench. In Section 4.2, we evaluate three SOTA NL-to-LTL frameworks on lifted NL to lifted TL translation. Note here, that measuring lifted translation performance on existing datasets is particularly difficult, as they present varying degrees of clarity in their lifted natural language elements. The CW [18], GLTL [10], and Navi [27] datasets have been processed to include lifted natural language components by Chen et al. [6] - and we perform similar processing of the Conformal dataset [28] to include it in our evaluation. In Section 4.3 we develop and evaluate two grounding baselines on our three datasets. In Section 4.4, we assemble the best results from the three individual evaluations to perform the first end-to-end translation evaluation. In Section 4.5 we perform our novel verification evaluation over the example traces of our dataset.
@@ 4.1 Lifting Evaluation
First, we evaluate four language models on the natural language lifting task. The LLM-based approaches each use the lifting prompt template from the NL2TL framework [6], which includes few-shot ground-truth NL to lifted NL examples from each of the datasets. The input to both models is a natural-language sentence and we compare the prediction made by the model against the ground-truth lifted natural language and measure the average IoU scores over every individual NL AP. We present the results in Table 2 where we see the linguistic complexity of our datasets is highlighted in the accuracies as even the best scoring model (GPT-4.1) reduces in performance on our new datasets. While this performance drop is minor on this model, it is more significant on the lower-cost, smaller GPT models. Regardless, this indicates our success in increasing the evaluation complexity.
Table 2: Comparison of lifting approaches. This table displays the mean IoU between predicted lifted natural language APs and ground-truth lifted natural language APs.
Accuracy (%)
S&R Traffic Light Warehouse
Model GLTL [10] CW [18] Conformal [28] Navi [27] (ours) (ours) (ours)
@@ 4.2 Lifted Translation Evaluation
Next, we evaluate the lifted translation capabilities of the three NL-to-LTL frameworks - nl2spec, NL2LTL, and NL2TL. In order to analyze the performance of their lifted translation abilities, the ground-truth lifted NL specification is given to the translation model, and the resulting lifted LTL translation is compared against the ground-truth lifted LTL. We perform this evaluation as a binary semantic accuracy using the pyModelChecking library [5] and present the results in Table 3. Here, we see that lifted translation can be very successful with both out-of-the-box LLM prompting (nl2spec) and with fine-tuned seq2seq models. However, as we have noted, we will see in end-to-end evaluation that this is an overconfident estimation of translation performance as grounding is not considered.
Table 3: Binary Accuracy on the Lifted NL to Lifted LTL task across seven datasets, for three frameworks (NL2LTL, nl2spec, NL2TL) given ground-truth lifted NL specifications.
Accuracy (%)
S&R Traffic Light Warehouse
Framework Translation Model GLTL [10] CW [18] Conformal [28] Navi [27] (ours) (ours) (ours)
@@ 4.3 Grounding Evaluation
In this section, we present the results obtained from our evaluation of our baseline grounding framework, applied to the ground truth lifted TL from our three VLTL-Bench datasets. We use two prompting strategies (described in Appendix A.4) applied to three GPT models to provide a broad evaluation of current grounding capabilities. Our first prompting baseline - few-shot - is composed of a brief description of the task at hand, accompanied by nine few-shot examples of correct (sentence, lifted sentence, AP-dictionary) tuples from all three scenarios (as opposed to individual scenarios). The next strategy is the scenario baseline prompt which includes the full scenario configuration file, as well as three few-shot examples from the dataset. To measure grounding accuracy, we parse the resulting AP-dictionary predictions and compare them with our ground-truth knowledge of the AP-dictionary in each entry. Our metrics are per-AP and per-AP-dictionary accuracy. Per-AP accuracy is calculated by recording the total number of correctly grounded APs divided by the total number of APs in the test set, and per-AP-dictionary accuracy is calculated by recording the total number of correct complete AP-dictionaries, divided by the size of the test set. These results are presented in Table 4.
@@
Table 4: Comparison of Grounding approaches. This table displays binary accuracy between predicted AP Grounding and known AP dictionary. LLM Baseline uses 9 few–shot sentence + lifted sentence + AP dict examples from every dataset; “Scenario” includes the scenario definition in the prompt and 3 examples from only that dataset.
Accuracy (of APs) Accuracy (of AP Dictionaries)
@@ 4.5 Verification Evaluation
NL2LTL [8] Few-shot 61.6 61.4 35.4 64.6 60.2 38.4 52.4 58.6 26.2
Scenario 1.06 32.0 7.4 61.8 59.2 36.6 12.4 36.2 9.8
nl2spec [7] Few-shot 47.4 48.0 34.8 47.2 46.0 33.6 46.0 44.2 29.6
Scenario 34.0 36.4 21.0 40.2 41.8 28.2 32.0 34.6 19.0
NL2TL [6] Few-shot 75.0 79.4 54.4 80.2 80.6 60.8 71.4 74.8 46.2
Scenario 27.5 50.8 22.1 72.6 76.3 54.5 33.3 52.4 23.5
@@ 5 Conclusion
We present the Verifiable Linear Temporal Logic Benchmark. VLTL-Bench is a suite of three new NL-to-LTL translation datasets that include the standard natural language and LTL pairs, supplemented with lifted natural language, lifted LTL, and trace examples. These additional features provide a method for the isolated training and evaluation of individual NL-to-LTL translation framework components. The provision of trace examples in VLTL-Bench introduces the possibility of a new type of input that is plausible in real-world translation frameworks, but unrepresented in current corpora. We acknowledge that the datasets included in the VLTL-Bench suite are generated using a finite number of linguistic and logical templates, populated by diverse synthetic natural language APs. Therefore, our logical diversity is limited to 43 lifted NL-LTL templates. The linguistic complexity of our benchmark is also partially limited in the same manner. However, our randomly selected action and target references combined with LLM-based pruning allows this work to surpass existing datasets in terms of diversity and realism, and we believe future effort should be partially devoted to further improvements in this area. VLTL-Bench reveals significant weaknesses in what were previously ironclad NL-to-LTL translation frameworks. Among these our the reliance on accurately lifted NL inputs for translation, lack of accurate grounding components, and lack of example trace inputs in current approaches. We hope that contribution will encourage exploration of diverse methods for NL-to-LTL translation, beyond the use of LLMs.
@@ Appendix A Appendix
In this appendix, we present a detailed overview of linear temporal logic in A.1, a discussion of verification via Kripke structures in A.2, a quantitative comparison of our VLTL-Bench dataset against existing datasets as well as examples from those datasets in A.3, our developed prompts for the baseline grounding approaches in A.4, the configuration files for our three scenarios in A.5, and finally our estimated compute resource usage and our external code and license information in A.6.
@@ A.1 Linear Temporal Logic
Linear temporal logic (LTL) is a modal extension of classical propositional logic that enables reasoning about how truths evolve over a discrete, linear timeline [32]. Formulas in LTL are interpreted over infinite sequences (or “traces”) of states
@@
Although our focus is on discrete-time LTL, many of these ideas carry over to related formalisms such as signal temporal logic (STL) for continuous-time, real-valued signals [19].
@@ A.2 Verification via Kripke Structures and Fluents
Verification of LTL specifications is typically conducted using a Kripke structure, which is a formal transition system comprising states, transitions, and labels indicating which atomic propositions hold true in each state. Formally, a Kripke structure is defined as a tuple $M=(S,S\_0,R,L)$, where:
@@
[…]
@@
A.3 Existing Datasets
Table 7: Comparison of NL–LTL datasets. Note that VLTL-Bench datasets include satisfying and unsatisfying traces for each entry. VLTL-Bench separates the total data into training and testing, as well as a subset of data taken from training and testing partitions that was selected by GPT-4o for its particularly realistic natural language.
Dataset Size Unique TL # APs # Actions # Args
GLTL Gopalan et al. [10] 11,109 37 4 1 4
CW MacGlashan et al. [18] 3,371 37 4 1 4
Conformal Wang et al. [28] 1,000 212 239 4 235
Navi Wang et al. [27] 7,474 6,414 221 26
Search-and-rescue [VLTL-Bench] 7,304 5,425 220 7 44
Traffic-light [VLTL-Bench] 7,319 6,196 5,046 4 175
Warehouse [VLTL-Bench] 7,457 5,991 5,074 5 452

### v2

@@ 3.3 Metrics
In this section, we introduce four complementary metrics that capture performance at different levels of the NL-to-TL pipeline, which is illustrated in Figure 2. Lifting accuracy measures the identification of atomic proposition spans in natural language, grounding accuracy evaluates their mapping to world state conditions, translation accuracy assesses logical equivalence between predicted and reference formulas, and verification accuracy checks whether predicted formulas satisfy or violate traces as expected. Together, these metrics provide a comprehensive view of system performance.
Lifting accuracy. For each token $\mathbb{S}_{i}$ in a sentence, the system predicts a label $\hat{\lambda}(\mathbb{S}_{i})\in\{0,1,\dots,k\}$, where 0 denotes background and $n$ denotes membership in $\pi_{n}$.
$\text{LiftAcc}\;=\;\frac{1}{|\mathbb{S}|}\sum_{i=1}^{|\mathbb{S}|}\bigl[\hat{\lambda}(\mathbb{S}_{i})=\lambda(\mathbb{S}_{i})\bigr].$
This measures the token-level classification accuracy of mapping substrings to atomic propositions.
Translation accuracy. Given a natural language specification $s$, the system produces a predicted TL formula $\hat{\varphi}$. Translation accuracy is an exact match between the predicted and reference formulas:
$\text{TransAcc}\;=\;\bigl[\hat{\varphi}\equiv\varphi\bigr],$
where $\equiv$ denotes logical equivalence. When working with lifted NL, the target is $\varphi_{L}$; for grounded NL, the target is $\varphi_{G}$.
Grounding accuracy. Let $\{\texttt{prop}_{1},\dots,\texttt{prop}_{k}\}$ be lifted placeholders and $g_{\mathcal{S}}$ the gold grounding function. The system predicts $\hat{g}_{\mathcal{S}}$.
$\text{GroundAcc}\;=\;\frac{1}{k}\sum_{j=1}^{k}\bigl[\hat{g}_{\mathcal{S}}(\texttt{prop}_{j})=g_{\mathcal{S}}(\texttt{prop}_{j})\bigr].$
This measures how well predicted atoms match their reference predicates and arguments.
Verification accuracy. For each dataset entry, two traces are provided: a positive trace $\sigma_{good}$ (satisfies $\varphi_{G}$) and a negative trace $\sigma_{bad}$ (violates $\varphi_{G}$). Given a predicted grounded formula $\hat{\varphi}_{G}$, verification checks whether the satisfaction relation holds:
$\text{VerifAcc}\;=\;\frac{1}{2}\Big([\sigma_{good}\models\hat{\varphi}_{G}]\;+\;[\sigma_{bad}\not\models\hat{\varphi}_{G}]\Big).$
@@ 3.4 Datasets
We construct four scenario definitions accompanied by action and target references, namely a Kitchen Assistant, Traffic Light, Search & Rescue, and Warehouse scenario. The details are provided in Appendix A.8. Using our proposed data synthesis, we generate four new datasets for training and evaluation. Each of our four datasets is designed to highlight distinct challenges for NL-to-LTL translation: the Traffic Light Control scenario is intended to balance action and argument grounding challenges, including a large library of “street name” arguments, but a smaller set of actions; the Search-and-Rescue scenario emphasizes multi-step temporal dependencies and deliberately includes ambiguous actions such as “avoid” and “communicate” to stress-test the system’s ability to distinguish between natural language verbs and temporal operators; the Kitchen Assistant scenario includes intentionally ambiguous action and argument references in order to stress both action and argument grounding simultaneously; and the Warehouse scenario introduces high semantic and linguistic variability by incorporating all 80 COCO object classes, making grounding especially complex. In this section, we use an entry from the Warehouse dataset as an example to illustrate the structure and properties of our data; additional examples from the other scenarios are provided in Appendix A.7.
Table 2: Comparison of NL–LTL datasets. We report the total number of entries (Size), the total number of unique TL entries, and the total number of unique APs appearing in the TL entries. Note that these datasets do not explicitly provide quantities of actions and arguments, and these are estimated by the authors.
Dataset Size Unique TL # APs # Actions # Args
GLTL GLTL-data 11,109 37 4 1 4
CW CW_data 3,371 37 4 1 4
Conformal conformalLTL_data 1,000 212 239 4 235
Navi Navi_data 7,474 6,414 221 26
Kitchen Assistant [VLTL-Bench] 10,000 7,385 172 9 36
Search-and-rescue [VLTL-Bench] 7,304 5,425 220 7 44
Traffic-light [VLTL-Bench] 7,319 6,196 5,046 4 175
Warehouse [VLTL-Bench] 7,457 5,991 5,074 5 82
@@
Our Warehouse dataset simulates a realistic warehouse retrieval scenario, explicitly designed for scalability and complexity in grounding tasks. Warehouse is our most distinct dataset with its inclusion of all 80 COCO (lin2014microsoft) object classes, significantly enriching the semantic and linguistic complexity and variation of atomic propositions. As with each of our datasets, all entries include LTL formulas with explicit grounding and alignment at token-level granularity, as well as verified positive (”good”) and negative (”bad”) execution traces for robust validation.
@@
  • Grounded LTL Formula: globally( deliver(bench, loading_dock) or idle() or search(wine_glass))
  • APs: prop_1 = “drop off long chair to loading dock”, prop_2 = “wait”, prop_3 = “look for glass for alcoholic beverage”
  • Positive Trace: [deliver(bench, loading_dock)], [idle()], [search(wine_glass)]
  • Negative Trace: [idle()], [idle()], [search(wine_glass), deliver(bench, loading_dock)]
@@ 4 Experimental Results
In this section, we present the results of multiple evaluations of NL-to-LTL translation frameworks and components. In Section 4.1, we measure the performance of common natural language lifting approaches, evaluated on four existing datasets in addition to three of the four datasets we present in VLTL-Bench. In Section 4.2, we evaluate three SOTA NL-to-LTL frameworks on lifted NL to lifted TL translation. Note here, that measuring lifted translation performance on existing datasets is particularly difficult, as they present varying degrees of clarity in their lifted natural language elements. In both translation evaluations, we use the pyModelChecking library (pymodelcheck_code) to determine logical equivalence. The CW (CW_data), GLTL (GLTL-data), and Navi (Navi_data) datasets have been processed to include lifted natural language components by (NL2TL), and we perform similar processing of the Conformal dataset (conformalLTL_data) to include it in our evaluation. In Section 4.3 we develop and evaluate two grounding baselines on our three datasets. In Section 4.4, we assemble the best results from the three individual evaluations to perform the first end-to-end translation evaluation. In Section 4.5 we perform our novel verification evaluation over the example traces of our dataset.
@@ 4.1 Lifting Evaluation
First, we evaluate four language models on the natural language lifting task. The LLM-based approaches each use the lifting prompt template from the NL2TL framework (NL2TL), which includes few-shot ground-truth NL to lifted NL examples from each of the datasets. The input to both models is a natural language sentence and we compare the prediction made by the model against the ground-truth lifted natural language using the lifting accuracy metric defined in Section 3.3. We present the results in Table 3 where we see the linguistic complexity of our datasets is highlighted in the accuracies, as even the best scoring model (GPT-4.1) reduces in performance on our new datasets. This performance drop is even more significant on the lower-cost, smaller GPT models. This indicates our success in increasing evaluation complexity.
Table 3: Comparison of lifting approaches.
Mean LiftAcc (%)
S&R TL WH
Model GLTL CW CF Navi (ours) (ours) (ours)
@@ 4.2 Lifted Translation Evaluation
Next, we evaluate the lifted translation capabilities of the three NL-to-LTL frameworks—nl2spec, NL2LTL, and NL2TL. In order to analyze the performance of their lifted translation abilities, the ground-truth lifted NL specification is given to the translation model, and the resulting lifted LTL translation is compared against the ground-truth lifted LTL. The formula for the translation accuracy metric is given in Section 3.3. We present these results in Table 4. Here, we see that lifted translation can be very successful with both out-of-the-box LLM prompting (nl2spec) and with fine-tuned seq2seq models. However, as we have noted, we will see in end-to-end evaluation that this is an overconfident estimation of translation performance as grounding is not considered.
Table 4: Comparison of four frameworks on the Lifted NL to Lifted TL translation task. Note that we provide ground-truth lifted NL specifications.
TransAcc (%)
S&R TL WH
Framework Model GLTL CW CF Navi (ours) (ours) (ours)
@@ 4.3 Grounding Evaluation
In this section, we present the results obtained from our evaluation of our baseline grounding framework, applied to the ground truth lifted TL from our three VLTL-Bench datasets. We use two prompting strategies (described in Appendix A.6) applied to three GPT models to provide a broad evaluation of current grounding capabilities. Our first prompting baseline—few-shot—is composed of a brief description of the task at hand, accompanied by nine few-shot examples of correct (sentence, lifted sentence, AP-dictionary) tuples from all three scenarios (as opposed to individual scenarios). The next strategy is the scenario baseline prompt which includes the full scenario configuration file, as well as three few-shot examples from the dataset. Our final grounding baseline employs the same scenario-specific few-shot examples as the previous approach, with the addition of specific instructions to include intermediate reasoning steps used to arrive at the answer. We use GPT-4o for this chain-of-thought approach in order to evaluate the performance of reasoning-capable models on this task. All four models are instructed to format their final answer in JSON format. To measure grounding accuracy, we parse the resulting AP-dictionary predictions and compare them with our ground-truth knowledge of the AP-dictionary in each entry. Our metrics are per-AP and per-AP-dictionary accuracy. Per-AP accuracy is calculated by recording the total number of correctly grounded APs divided by the total number of APs in the test set, and per-AP-dictionary accuracy is calculated by recording the total number of completely correct AP-dictionaries, divided by the size of the test set. These results are presented in Table 5.
@@
Table 5: Comparison of Grounding approaches. This table displays binary accuracy between predicted AP Grounding and known AP dictionary. LLM Baseline uses 9 few–shot sentence + lifted sentence + AP dict examples from every dataset; “Scenario” includes the scenario definition in the prompt and 3 examples from only that dataset. Note that Lang2LTL grounds using cosine similarity between reference and canonical AP embeddings.
Accuracy (% of APs) Accuracy (% of AP Dictionaries)
@@ 4.5 Verification Evaluation
NL2LTL (NL2LTL) Few-shot General 61.6 61.4 35.4 64.6 60.2 38.4 52.4 58.6 26.2
Few-shot Scenario 1.06 32.0 7.4 61.8 59.2 36.6 12.4 36.2 9.8
nl2spec (nl2spec) Few-shot General 47.4 48.0 34.8 47.2 46.0 33.6 46.0 44.2 29.6
Few-shot Scenario 34.0 36.4 21.0 40.2 41.8 28.2 32.0 34.6 19.0
NL2TL (NL2TL) Few-shot General 75.0 79.4 54.4 80.2 80.6 60.8 71.4 74.8 46.2
Few-shot Scenario 27.5 50.8 22.1 72.6 76.3 54.5 33.3 52.4 23.5
Lang2LTL (Lang2LTL) Embedding 43.3 61.9 39.3 44.7 63.0 41.3 21.6 40.1 16.6
@@ 5 Conclusion
We present the Verifiable Linear Temporal Logic Benchmark. VLTL-Bench is a suite of three new NL-to-LTL translation datasets that include the standard natural language and LTL pairs, supplemented with lifted natural language, lifted LTL, and trace examples. These additional features provide a method for the isolated training and evaluation of individual NL-to-LTL translation framework components. The provision of trace examples in VLTL-Bench introduces the possibility of a new type of input that is plausible in real-world translation frameworks, but unrepresented in current corpora. We acknowledge that the datasets included in the VLTL-Bench suite are generated using a finite number of linguistic and logical templates, populated by diverse synthetic natural language APs. VLTL-Bench reveals significant weaknesses in what were previously ironclad NL-to-LTL translation frameworks. Among these weakness are: the reliance on accurately lifted NL inputs for translation, lack of accurate grounding components, and lack of example trace inputs in current approaches. We envision our contribution will encourage exploration of diverse methods for grounded NL-to-LTL translation, beyond the use of LLMs.
@@ Appendix A Appendix
In this appendix, we present a detailed overview of linear temporal logic in A.1, a discussion of verification via Kripke structures in A.2, a quantitative comparison of our VLTL-Bench dataset against existing datasets as well as examples from those datasets in A.5, our developed prompts for the baseline grounding approaches in A.6, the configuration files for our three scenarios in A.8, and finally our estimated compute resource usage and our external code and license information in A.9.
@@ A.1 Linear Temporal Logic
Linear temporal logic (LTL) is a modal extension of classical propositional logic that enables reasoning about how truths evolve over a discrete, linear timeline (LTL). Formulas in LTL are interpreted over infinite sequences (or “traces”) of states
@@
Although our focus is on discrete-time LTL, many of these ideas carry over to related formalisms such as signal temporal logic (STL) for continuous-time, real-valued signals (STL).
@@ A.2 Verification via Kripke Structures and Fluents
Verification of LTL specifications is typically conducted using a Kripke structure, which is a formal transition system comprising states, transitions, and labels indicating which atomic propositions hold true in each state. Formally, a Kripke structure is defined as a tuple M=(S,S_0,R,L)M=(S,S\_0,R,L), where:
@@
  • $L:S\rightarrow 2^{\mathit{AP}}$ is a labeling function mapping states to the sets of atomic propositions that are true in each state.
@@
A.3 VLTL-Bench LTL Expression Statistics
Token / Operator Search & Rescue Traffic Light Warehouse
and 5536 5508 5297
double_implies 1104 1141 1091
finally 3835 3842 3918
globally 9899 9851 9820
implies 5164 5295 5264
next 12144 12304 11713
not 6781 6724 6593
or 3198 3229 3054
prop_1 14440 14466 14193
prop_2 7934 7964 7835
prop_3 3740 3831 3825
until 1112 1088 1147
Table 8: Operator splits and template breakdowns by domain.
A.4 VLTL-Bench New Templates
We then craft 7 of our own templates to fill perceived gaps in specification coverage. Of these templates, 4 entries include new lifted LTL halves (marked below with a *), and 3 include new lifted NL halves.
NL LTL
finally ( not prop_1) “eventually, avoid prop_1”
globally ( not prop_1) “always avoid prop_1”; “prop_1 must never occur”
next prop_1 “at the next time step, prop_1 holds”
prop_1 until prop_2 “prop_1 must always hold at all times before prop_2”
finally (prop_1 and prop_2) OLD: “Eventually, both prop_1 and prop_2 will hold simultaneously”
NEW: “At some point, prop_1 and prop_2 will both hold at the same time.”
globally (prop_1 and prop_2) OLD: “Both prop_1 and prop_2 hold at every step.”
NEW: “At all time steps, prop_1 and prop_2 both hold.”
finally (prop_1 or prop_2) OLD: “eventually, either prop_1 or prop_2”
NEW: “either prop_1 or prop_2 will hold at some point in time.”
Table 9: Examples of NL–LTL mappings. OLD/NEW entries show updated phrasing.
A.5 Existing Datasets